# Neural models of multiscale systems: conceptual limitations, stochastic parametrizations, and a climate application

### v2

@@ IV.2.4 Main takeaways
In this Section, we developed a stochastic neural emulator that captures the joint evolution of the sea surface temperature (SST) field and the global mean net radiative flux at the top of the atmosphere (TOA) in the GFDL-CM4 climate model. Addressing a specific, well-posed problem, proved crucial in guiding the construction of the model: it enabled us to build on previous literature, restrict attention to just two variables and a limited range of spatial and temporal scales. These simplifications led to a stable and computationally efficient emulator that responds to external forcing and can be used to investigate causal relationships via perturbation experiments. The model reconstruct the net TOA radiative flux when forced with a pre-computed SST trajectory. Beyond reconstruction, we illustrated that the model can be used to perform perturbation experiment, allowing us to enhance mechanistic understanding of the underlying physical processes. Findings are in qualitative agreement with results from traditional perturbation experiments in numerical models, giving additional confidence in the emulator’s ability to capture key interactions across variables.
V Conclusions
In the first part of this paper, we addressed, at a pedagogical level, fundamental issues in neural emulation of dynamical systems with climate applications in mind. We considered two scenarios: (i) when the full state vector is observed, and (ii) the more realistic case of partial observations. Our findings can be summarized as follows:
@@ V Conclusions
    Fully observed state vector. In this idealized case, it is possible to infer a dynamical model from data alone that reproduces the system’s stationary statistics and responds accurately to small, external perturbations.
@@
    Partially observed state vector. This more realistic case presents fundamental difficulties, even with access to high-quality data and powerful neural models architectures. The model skill now depends critically on two factors: (a) the choice of variables to model and (b) how unobserved degrees of freedom are parameterized.
These results highlight intrinsic limitations of purely data-driven modeling in high-dimensional systems, where the full set of relevant variables is typically unknown and there is no rigorous mathematical framework for selecting the proper variables to model. The identification of such variables, when possible, is inherently problem-dependent and often leads to adopting a coarse-grained representation of the dynamics. The next challenge is to parameterize the cumulative effect of unobserved degrees of freedom. This final step underscores that the issue of parameterization, traditionally associated with numerical modeling, is equally central to data-driven approaches using neural emulators. A promising perspective is to reconcile neural networks with principled modeling strategies, such as reduced-order modeling and stochastic parameterizations, by focusing on coarse-grained dynamics rather than constructing deterministic models that attempt to resolve as many spatiotemporal scales as possible.
@@
In addition, we argued that impulse response functions are a building block that a model should capture in order to reproduce causal interactions across degrees of freedom and to accurately respond to external perturbations. This perspective enables us to move beyond evaluation metrics based solely on stationary statistics and instead consider the effect of external interventions. Assessing impulse responses is straightforward in the case of data-driven emulators built from simple numerical models, i.e. models that allow us to simulate thousands of ensemble members, as in the first part of this study. In such cases, one can actively perturb the reference model to generate a ground truth for the response operator. However, for models constructed from observational data, such direct perturbations are not possible. In this context, one could, in principle, leverage the Fluctuation-Dissipation Theorem (FDT) to estimate impulse response functions from data alone, at least at a qualitative level [38, 37, 35, 47, 62]. In particular, it has been shown that even simple approximations of FDT, such as the quasi-Gaussian approximation [37], often yield qualitatively reliable estimates of impulse responses in the ensemble mean, without the need to perturb the system directly; see, for example, [122, 123, 104]. That said, FDT itself is inferred from data and should therefore be viewed as a complementary, theory-based diagnostic (rather than a ground truth) allowing us to cross-validate the inferred responses of the emulator. In other words, consistency between the responses predicted by the emulator and those estimated via FDT can increase confidence in the model’s ability to capture the system’s response in the ensemble mean to external perturbations. In contrast, assessing impulse responses of higher-order moments (e.g., variance) via FDT remains substantially more challenging and calls for new methodological tools. A promising ongoing direction involves score-based generative modeling, as explored in [124, 125]. Another fruitful avenue lies in data-driven frameworks at the intersection of Markov State Modeling and Koopman operator theory [126, 127]. It is worth emphasizing that the central limitation discussed earlier, i.e. the need to choose the proper variables, remains relevant for the inference of response operators via FDT as well (and for any causal inference methodology); see the discussion in [47].
Building on the theoretical insights introduced in the first part, the second part of the paper presents a stochastic neural emulator for the coarse-grained dynamics of sea surface temperature (SST) and net top-of-atmosphere (TOA) radiative flux. The model was trained on long unforced simulations of a coupled climate model and designed with a specific physical task in mind: capturing radiative feedbacks to SST anomalies. We showed that the emulator reproduces the stationary statistics of the system and, more importantly, allows us to reconstruct TOA radiative fluxes in forced simulations when provided with forced SST trajectories. To deepen this analysis, we applied linear response theory to extract the direct causal links between temperature modes and TOA fluxes, targeting the so-called “pattern effect” [30, 29]. These links, when projected back onto the high-dimensional state space, were found to be consistent with recent studies. This suggests that the emulator’s skill in reconstructing TOA fluxes under forcing arises for the right physical reasons. In this way, focusing on a few dominant modes, on a limited range of time scales, and using stochastic parametrizations allowed us to build a skillful and interpretable emulator.
We emphasize that the fundamental limitations discussed here apply to data-driven models of complex dynamical systems that must both represent stationary statistics and respond faithfully to external perturbations. For Earth system models specifically, these considerations primarily concern climate emulators (particularly coupled atmosphere-ocean systems), and should not be conflated with the different challenge of weather forecasting using neural networks (e.g., [128]). Weather prediction focuses on short-term predictability where long-term stability and forced responses are of secondary importance.
Modeling real-world dynamical systems from data alone is severely constrained by high dimensionality and the fact that we typically observe only a subset of relevant variables [46, 32, 40]. This leads to the fundamental challenge of identifying, when possible, the proper variables to model, a problem recognized since the work of Onsager and Machlup (1953) [68] and highlighted in more recent studies [47, 48]. These challenges are especially pronounced in climate dynamics, where observations are always partial, and the underlying system is both high-dimensional and multiscale. In this context, the multiscale structure of the climate system can be an advantage, allowing us to model coarse-grained dynamics while treating unresolved processes stochastically. The need to formulate models tackling the effective dynamics of slow variables, rather than the full set of primitive equations, has a long history in climate science, even in non-data-driven settings where all governing equations are known. The early work of Charney and von Neumann, for example, demonstrated the conceptual and practical benefits of filtering out “meteorologically insignificant noise” and emphasized the need for interpretable and computationally efficient effective equations in metereology [129, 63, 43, 44]. These insights remain conceptually relevant, especially the focus on effective dynamics, in the development of neural emulators for climate, including coupled atmosphere–ocean systems. In purely data-driven settings, focusing on stochastic models that target coarse-grained processes offers a promising way forward by reconciling rigorous approaches such as reduced-order modeling with the flexibility and power of neural networks. Possible strategies include learning stochastic parameterizations directly from data using deep learning (as in this paper), guiding variable selection through causal inference [130, 131], incorporating information theory and response theory into the training phase [56], studying the evolution of the climate system in its “latent” dynamics [65] (i.e. on the attractor) and incorporating empirical knowledge and physics constraints in the learning phase [11, 7, 132, 25]. Finally, the rather hard difficulty (at times irreducible) of data-driven modeling for the coupled atmosphere–ocean system can at times be ameliorated by building models tailored to solve specific problems, by building on prior literature and insights, and developed with a clear understanding of the relevant variables and the spatiotemporal scales at play.
Appendix A Technical and practical details on the model formulations
A.1 Multivariate data-driven stochastic models

### v3

@@ IV.2.4 Main takeaways
In this Section, we developed a stochastic neural emulator that captures the joint evolution of the sea surface temperature (SST) field and the global mean net radiative flux at the top of the atmosphere (TOA) in the GFDL-CM4 climate model. Addressing a specific, well-posed problem, proved crucial in guiding the construction of the model: it enabled us to build on previous literature, restrict attention to just two variables and a limited range of spatial and temporal scales. These simplifications led to a stable and computationally efficient emulator that responds to external forcing and can be used to investigate causal relationships via perturbation experiments. Such experiments focused on very large ensembles ($10^{6}$ trajectories), currently impossible to perform with climate models. These results demonstrate that reduced-order neural emulators can extend beyond reproducing stationary climate dynamics to enable efficient, ensemble-based perturbation experiments.
V Discussion and Conclusion
In summary, the first part of this work discussed the limitations of purely data-driven models of complex climate systems. We highlighted the potential of combining traditional reduced-order modeling approaches, including state-dependent (multiplicative) stochastic parameterizations, with recent advances in neural networks to study large-scale, low-frequency climate dynamics. We emphasized the need to evaluate such models against both stationary and perturbed statistics. In this context we discussed how response theory provides rigorous guidance to move beyond stationary evaluation and probe causal mechanisms. These ideas were then illustrated through a climate-dynamics problem focusing on inferring causal relations between surface warming and changes in net radiative fluxes. Given careful variable selection, reduced-order neural stochastic models can capture both stationary and perturbed statistics in a computationally efficient way. Below, we expand on these points and outline directions for future work.
V.1 Limits of data-driven-only climate emulators and the role of stochastic approaches
In the first part of this paper, we distinguished two possible scenarios in data-driven modeling:
@@ V.1 Limits of data-driven-only climate emulators and the role of stochastic approaches
    Unknown equations, fully observed state vector. In this idealized setting, it is possible to infer a model from data alone that reproduces the system’s stationary statistics and responds accurately to small external perturbations.
@@
    Unknown equations, partially observed state vector. This more realistic case presents fundamental difficulties, even with access to high-quality data and powerful neural models architectures. The model skill now depends critically on two factors: (a) the choice of variables to model and (b) how unobserved degrees of freedom are parameterized.
These results highlight intrinsic limitations of purely data-driven-only modeling in high-dimensional climate systems, where the full set of relevant variables is typically unknown and no rigorous framework exists for selecting the proper variables to model. By variable selection we mean the choice of climate fields as well as their spatial and temporal scales (also including dimensionality reduction methodologies). Such choices are guided by the physical process of interest, and in the case of low-frequency climate dynamics, they motivate the development of reduced-order neural emulators tailored to specific phenomena as efficient alternative to general purpose emulators. A central challenge is to parameterize the cumulative effect of unresolved degrees of freedom. Model performance (especially in capturing responses to external forcings) depends critically on this choice and we showed that multiplicative (state-dependent) noise is preferable to additive noise.
We emphasize that the analysis presented here applies to data-driven models of turbulent geophysical systems that must capture both stationary statistics and responses to external perturbations, as in (causal) Green’s function experiments [38]. For Earth system models, these issues mainly concern autoregressive climate emulators, particularly coupled atmosphere–ocean systems, and should not be confused with short-term weather forecasting using neural networks (e.g., [115, 23]), where long-term stability and forced responses are less relevant. Our study also does not address statistics-based emulators for climate change projections, which have proven effective for investigating responses of key climate variables to CO2 forcing [116, 117]; see [116] for an example where simple approaches, such as linear pattern scaling, can outperform more complex models.
Future work.
The discussion above suggests two possible directions for future research in autoregressive emulators of climate. First, a stochastic formulation of neural autoregressive climate emulators is preferable to a purely deterministic one. Probabilistic formulations align more closely with the inherently stochastic and uncertain nature of climate dynamics [118]. Recent advances in weather emulation already point in this direction such as FourecastNet3 [23], GenCast [23] and the novel work of Alet et al. (2025) [119]. Second, the ability of climate emulators in capturing forced responses in the full probability distribution, akin to Green’s function experiments, remains largely unexplored. A promising way forward is to evaluate novel architectures on their ability to respond to external perturbations on a hierarchy of simplified models relevant to geophysical fluid dynamics, such as the Lorenz–96 system [120]. These controlled testbeds enable rigorous quantification of forced responses and could inform the design of future architectures for data-driven climate modeling.
V.2 Response theory as a theoretical framework for model evaluation
@@
We argued that the impulse response operator is a building block that a model should capture in order to reproduce causal interactions across degrees of freedom and to accurately respond to external perturbations (Section II.2 and Appendix A). This perspective enables us to move beyond evaluation metrics based solely on stationary statistics and instead consider the effect of external interventions. Assessing impulse responses is straightforward in the case of data-driven emulators built from simple numerical models, i.e. models that allow us to simulate thousands of ensemble members, as in the first part of this study. In such cases, one can actively perturb the reference model to generate a ground truth for the response operator. However, for models constructed from observational data, such direct perturbations are not possible. In this context, one could, in principle, leverage the Fluctuation-Dissipation Theorem (FDT) to predict the response operator from data alone [81, 17]. Traditionally, the quasi-Gaussian approximation have yield reliable estimates of impulse responses in the ensemble mean [121, 47]. Recent advances in score-based generative models have shown the potential to extend these predictions to responses in higher-order moments [114].
Future work.
In the absence of direct perturbation experiments, FDT provides a rigorous bridge between stationary and forced dynamics [49, 48]. With sufficiently long integrations of current climate emulators, one can readily focus on representative large-scale modes and estimate their impulse responses from the stationary record, thereby extending evaluation beyond stationary statistics with minimal computational cost. A recent example illustrating this approach in the climate context can be found in [18]. Moreover, the response operator at the shortest time step encodes direct causal links, i.e. the input–output relationships across variables in a reduced-order emulator (see Appendix A). Future work will leverage this information to suppress spurious, non-causal input-output dependencies during training, similarly to the recent work of Chen and Liu (2024) [59] on causally constrained, sparse reduced-order models.
V.3 A realistic application: radiative responses to temperature perturbations
Building on the theoretical insights introduced in the first part, the second part of the paper presented a stochastic neural emulator for the coarse-grained dynamics of sea surface temperature (SST) and net radiative flux at the top-of-atmosphere (TOA). The model was trained on a long unforced simulations of a coupled climate model and designed with a specific physical task in mind: capturing radiative feedbacks to SST anomalies. The emulator focuses on the evolution of 20 principal components of the SST field and on the scalar global mean, net radiative flux; unobserved degrees of freedom are parametrized through multiplicative noise through an additional neural network. The emulator largely reproduces the stationary statistics of the system and, more importantly, allows us to reconstruct TOA radiative fluxes in two forced simulations. We inferred the direct causal links between temperature modes and TOA fluxes through a perturbation approach, akin to Green’s function-like experiments with climate models [38], targeting the so-called “pattern effect” [39]. Crucially, the analyses targeted both responses in the ensemble mean and variance. Such responses are then projected back onto the EOFs to obtain maps of sensitivity in mean and variance change of the radiative flux at the TOA to SST warming perturbations. We argued that the sensitivity map for the ensemble mean is consistent with recent studies, indicating that the emulator’s skill in reconstructing TOA fluxes under forcing arises from physically meaningful mechanisms. In contrast, the sensitivity map of the ensemble variance is a novel result, not previously examined in the literature. It highlights the potential for uncertainty quantification in perturbation experiments but it requires further investigation to better understand the underlying physical mechanisms. These findings highlight the potential of reduced-order neural emulators to perform thousands of perturbation experiments, shifting the focus of climate feedback analysis from mean responses to the responses of probability distributions.
Modeling the climate system from data alone is constrained by observing only a subset of variables of a multiscale, high-dimensional dynamical system. This raises the fundamental challenge of identifying the appropriate variables (both fields and spatiotemporal scales) a fundamental problem in physics, emphasized in classical work [122] and revisited in recent studies [123]. However, large-scale climate dynamics is often well characterized by a few recurrent spatiotemporal patterns, or modes of variability. In this case, the multiscale structure of the climate system can be an advantage allowing us to derive data-driven models of the coarse-grained dynamics while treating unresolved, fast processes stochastically. Traditional reduced-order models stand out from general-purpose emulators by targeting specific climate problems from the outset, guided by prior literature, physical insight, and a clear understanding of the relevant variables and scales at play. Such data-driven modeling approach has a long history in climate science [2], though most efforts have emphasized reproducing stationary statistics or short-term forecasting [3, 4, 75] rather than forced responses. A promising perspective is to build on such established reduced-order approaches, focusing on powerful neural networks to fit deterministic and stochastic components and on tools from response theory, such as the FDT, to guide model development. Together, these strategies can enable efficient algorithms to study how full probability distributions respond to external perturbations, opening new avenues for causal inference and attribution in climate science.
Appendix A Response operator: practical estimation and causal interpretation
A.1 Inferring the response operator in practice
  1. 1.
    We simulate a very long trajectory of the model, starting from a random initial condition and remove an initial transient. We then sample $N_{e}$ random points from the simulation to define an ensemble of $N_{e}$ initial conditions on the model’s attractor. $N_{e}$ should be very large (i.e. $N_{e}\gg 1$) in order to sample the whole attractor and approximate averages over the invariant distribution.
  2. 2.
    For each one of the $N_{e}$ initial conditions, we impose an impulse perturbation $x_{j}(0)\rightarrow x_{j}(0)+\Delta_{j}\delta(t)$ of size $\Delta_{j}$ on the $j$-th degree of freedom at time $t=0$. $\delta(t)$ is the Dirac delta function. The amplitude $\Delta_{j}$ should be theoretically infinitesimally small, ensuring linearity of the response even in nonlinear systems. In practice we do as follows: we consider the long time series of $x_{j}(t)$ from the long control integration above and define $\Delta_{j}=10^{-2}\sigma_{j}$, where $\sigma_{j}$ is the standard deviation of time series $x_{j}(t)$.
  3. 3.
  4. 4.
    At each time $t$ we then estimate the time dependent ensemble average $\langle A(x_{k}(t))\rangle$ of an observable $A(x_{k}(t))$ for both the perturbed and unperturbed run. We refer to the perturbed and control ensemble average respectively as $\langle A(x_{k}(t))\rangle_{\text{p}}$ and $\langle A(x_{k}(t))\rangle$. The observables considered in this study are going to be: (i) $A(x_{k}(t))=x_{k}(t)$ and (ii) $A(x_{k}(t))=(x_{k}(t)-\mu_{k}(t))^{2}$, with $\bm{\mu}(t)$ representing the time-dependent mean of the distribution, therefore quantifying the response in ensemble mean and variance.
  5. 5.
    We define the impulse response operator for observable $A(x_{k}(t))$ as
    $R_{k,j}(t)=\frac{\langle A(x_{k}(t))\rangle_{\text{p}}-\langle A(x_{k}(t))\rangle}{\Delta_{j}}.$
A.2 Causal interpretation of the response operator
Causality in natural systems is inferred through external interventions: an agent perturbs the system and the response is measured [77, 78]. Climate modeling follows the same principle through Green’s function experiments, where one variable is perturbed and the response of another is evaluated [38, 124]. Linear response theory formalizes this principle, providing a rigorous framework for causal analysis both in models and directly from data [17, 18, 114].
Because responses are inherently time-dependent, causal links through responses can be distinguished in three forms, all encoded in the time-dependent response operator: (i) direct links, (ii) indirect links, and (iii) cumulative links (or causal sensitivity). In what follows, we use the triad system in Eq. (6) as a simple example to illustrate these ideas in a clear and unambiguous way. Finally, we summarize recent approaches for estimating responses directly from data and refer to Baldovin et al. (2020) [17] for details on the theoretical foundations of this approach.
Consider a $N$ dimensional dynamical system of the kind:
$\begin{split}\frac{dx_{i}}{dt}=f_{i}(x_{1},x_{2},...,x_{N});~\text{with}~i=1,2,...,N.\end{split}$ (14)
For stochastic systems, causal links are contained in the deterministic drift $f_{i}$, so the discussion below applies to both deterministic and stochastic dynamics. In this setting, the existence of a direct causal link $x_{j}\rightarrow x_{k}$ is determined by whether $x_{j}$ appears in the arguments of $f_{k}$. More formally, the absence of a direct causal link $x_{j}\rightarrow x_{k}$ is expressed by:
$\frac{\partial f_{k}}{\partial x_{j}}=0.$
As a concrete example, we return to the triad system in Eq. (6), for which the underlying direct causal links are known exactly. This example provides a straightforward, intuitive illustration of how perturbation experiments reveal causal interactions. Figure 8 reports both (a) its functional form and (b) its causal graph, which encode the same information.
$\frac{dx_{1}}{dt}$ $=f_{1}(x_{1},x_{2},x_{3})$
$\frac{dx_{2}}{dt}$ $=f_{2}(x_{1},x_{2})$
$\frac{dx_{3}}{dt}$ $=f_{3}(x_{1},x_{3})$
(a) Functional form of the triad system in Eq. (6).
(b) Causal graph of the triad system.
Figure 8: Panel (a) Functional form of the triad system in Eq. (6). Panel (b) Graph structure of the triad system. Note from panel (b) that causality in dynamical systems can involve cycles, in contrast with the acyclic structures often assumed in applications outside physics, see e.g. [125].
Causal relations can then be probed by adding independent impulse perturbations to each degree of freedom and record the system’s response. Responses are computed up to a long time horizon $\tau_{\max}$, with discrete times $t=0,dt,2dt,\dots,\tau_{max}$. Since we are interested in causal relations, it suffices to consider the identity observable $A(x_{k}(t))=x_{k}(t)$, i.e. the response of the ensemble mean [17]. We now clarify the three different types of causal links mentioned at the start of this Section.
Direct causal links.
The direct causal links are encoded in the response operator $\mathbf{R}(dt)$ evaluated at the shortest time step $t=dt$. The intuition follows directly from the graph representation in Fig. 8(b) and is rather straightforward.
  • Any perturbation applied to $x_{2}$ will affect $x_{2}$ itself and $x_{1}$ after one time step, but not $x_{3}$. Hence, we infer $\partial f_{3}/\partial x_{2}=0$.
  • A perturbation applied to $x_{3}$ will affect $x_{3}$ and $x_{1}$, but not $x_{2}$, so $\partial f_{2}/\partial x_{3}=0$.
  • By contrast, a perturbation to $x_{1}$ propagates to all three variables within a single step, indicating direct causal links from $x_{1}$ to both $x_{2}$ and $x_{3}$.
Indirect causal links.
Causal relations in physical systems extend beyond direct interactions. A variable $x_{j}$ may affect another variable $x_{k}$ indirectly through an intermediate variable $x_{i}$, forming a pathway $x_{j}\rightarrow x_{i}\rightarrow x_{k}$. Such indirect links are central in climate dynamics [126] and in complex systems more broadly [127]. Within the response framework, these pathways are revealed by the time-dependent operator $R_{k,j}(t)$: the first non-zero response identifies the shortest causal path in the graph connecting $x_{j}$ to $x_{k}$. Crucially, if $R_{k,j}(t)=0$ for all time $t$, then no external intervention on $x_{j}$ can ever influence the behavior of $x_{k}$.
Cumulative causal links or causal sensitivity.
Finally, Baldovin et al. [17] introduced a cumulative (in time) measure of causation between two variables, defined as $D_{k,j}=\int_{0}^{\tau_{max}}R_{k,j}(t)dt$. Such cumulative link, can be also referred to as “causal sensitivity”. Such a measure is especially relevant in practical experiments with high-dimensional models where interactions unfold across multiple spatial and temporal scales. A concrete example using climate models, is provided by the approach of Bloch-Johnson et al. (2024) [38], where causal links are inferred between patterns of surface warming and the net radiative flux at the top of the atmosphere. Their method reconstructs responses to perturbations by: (i) applying surface warming perturbations, (ii) integrating an atmospheric-only model with and without the perturbation, and (iii) computing the average (in time) difference between the two runs after removing an initial transient. In this case, the causal link emerges precisely as a cumulative (in time) atmospheric response to the surface perturbations.
We emphasize the crucial role of ensemble averages in establishing the general validity of response theory. Evaluating responses along single trajectories of nonlinear systems is strongly affected by chaos, restricting the applicability of these ideas to short times and very small perturbations. In contrast, focusing on observables, i.e. quantities averaged over the invariant density, ensures the robustness of the theory and largely removes the dependence on chaos. For a detailed discussion of this point, we refer the reader to [128]. Future efforts to estimate responses in climate science, which often rely on perturbing single trajectories, would therefore benefit from focusing on ensemble-averaged observables.
Response operator from data alone.
Given the assumption of Markovianity it is possible to infer the system’s response operator from a long stationary trajectory using the Fluctuation-Dissipation Theorem (FDT) [17]. The FDT formula provides a link between stationary and perturbed dynamics by predicting forced responses from stationary statistics alone [129]. Specifically:
$R_{k,j}(t)=-\Big\langle A(x_{k}(t))\frac{\partial\ln\rho(\mathbf{x})}{\partial x_{j}}\Big|_{\mathbf{x}(0)}\Big\rangle,$ (15)
where $A(x_{k}(t))$ is a general observable, $\rho(\mathbf{x})$ is the stationary distribution. $\langle\cdot\rangle$ refers to ensemble average, approximated through temporal averages. $\frac{\partial\ln\rho(\mathbf{x})}{\partial x_{j}}\Big|_{\mathbf{x}(0)}$ represents the so-called “score function” evaluated at the initial state, i.e. on the attractor [15]. Therefore, in principle the response operator can be computed from data if the score $\nabla\ln\rho(\mathbf{x})$ is available. In practice, the score is often inaccessible, and most applications rely on the quasi-Gaussian (qG) approximation, by replacing $\rho(\mathbf{x})$ with a Gaussian distribution $\rho^{G}(\mathbf{x})$ with same mean and covariances [130, 48]. For responses in the ensemble mean—which are sufficient to infer causal links—the qG approximation achieves high skill even for nonlinear systems [121, 17, 114]. Recent work [114] has demonstrated that generative modeling techniques can be used to efficiently estimate the score directly from data, leading to high skill in computation of higher-order responses (up to the 5th moment) in strongly nonlinear regimes. Other fruitful avenues lie in frameworks at the intersection of Markov State Modeling and Koopman operator theory such as [131].
It is worth emphasizing that the central limitation discussed in the main text, i.e. the need to choose the proper variables, remains relevant for the inference of response operators via FDT as well (and for any causal inference methodology); see the discussion in [17].
Appendix B Relevance of the the triad system for geophysical turbulence
Here, we motivate the choice of the triad model analyzed in Section III as a building block for geophysical turbulent flows. We further present and test the associated analytical stochastic model for the slow variable $x_{1}$ (as discussed in Section III.3.1), illustrating that Markovian reduced-order models can sometimes be derived from partial observations. For detailed mathematical derivations and additional examples, we refer to the original work by Majda and collaborators [46, 60, 61, 48, 16].
B.1 Link to geophysical turbulence
Projecting many geophysical turbulent flows, such as quasigeostrophic models, onto a complete set of orthogonal basis functions (spectral or empirical modes such as EOFs), give rise to the following generic form for the mode amplitudes:
$d\mathbf{u}/dt=\mathbf{F}+\mathbf{L}\mathbf{u}+\mathbf{B}(\mathbf{u},\mathbf{u}),$ (16)
where $\mathbf{F}$ is an external forcing, $\mathbf{L}$ is a linear operator and $\mathbf{B}(\mathbf{u},\mathbf{u})$ is a nonlinear, quadratic operator. The linear operator $\mathbf{L}=\mathbf{L}^{\prime}+\mathbf{D}$, is formed by a skew-symmetric operator $\mathbf{L}^{\prime}$ and by a diagonal dissipation matrix $\mathbf{D}$ with negative real entries. The quadratic nonlinearity should be energy-conserving, satisfying $\mathbf{u}\cdot\mathbf{B}(\mathbf{u},\mathbf{u})=0$ [132, 60, 48]. Eq. (16) is the typical abstract form of dry cores of atmospheric models [67]. Galerkin truncations of turbulent systems often reduce to this structure, as in the Lorenz ’63 and Charney–DeVore models [132, 133, 68]. $\mathbf{u}\in\mathbb{R}^{N}$ represents the full turbulent geophysical system, with $N$ infinite in principle and finite but very large in applications ($N\gg 1$). The state $\mathbf{u}$ is then decomposed into slow and fast components, $\mathbf{u}=(\mathbf{x},\mathbf{y})$, i.e. $\mathbf{x}$ will represent the first $n$ components of $\mathbf{u}$ and $\mathbf{y}$ the next NnN-n.
We consider the simple case of $N=3$: one large-scale resolved mode $x_{1}$ and two fast modes $\mathbf{y}=(x_{2},x_{3})$. Majda and collaborators [46] considered an exact projection of the slow mode $x_{1}$ and the two fast modes $(x_{2},x_{3})$ onto the dynamics in Eq. (16). This gives rise to three equations, defined by a forcing term, linear and quadratic self- and cross-interaction across slow and fast time scales. The triad model in Eq. (6) is then obtained by approximating the quadratic self-interactions of the fast modes as a damping and stochastic forcing, as: $\gamma_{i}\epsilon^{-1}x_{i}+\sigma_{i}\epsilon^{-1/2}\xi_{i}$, with $i=2,3$. $\xi_{i}$ is a standard Gaussian white noise process. $\epsilon$ sets the time-scale separation between the slow and fast modes. The key assumption is that the detailed nonlinear self-interactions among fast modes are not important if the goal is to capture the coarse-grained slow dynamics. This approximation transforms the original deterministic equations into a stochastic system. The specific version used in Eq. (6) is further simplified by setting to zero a few interactions, such as the linear cross- and self-interactions of the fast modes as well as the external forcing terms.
Therefore, the relevance of the triad system (Eq. (6)) to this study is that it provides a minimal multiscale setting with linear and quadratic nonlinear interaction of the kind that naturally arise when geophysical turbulent flows are expressed in terms of orthogonal bases such as EOFs. For additional details see also Section 4 of [16].
B.2 Analytical stochastic model for
the slow variable $x_{1}$
Starting from the triad model in Eq. (6), Majda and collaborators [60] showed that an analytical reduced-order equation for the slow variable $x_{1}$ can be rigorously derived in the limit $\epsilon\to 0$ using standard projection techniques [48, 46, 47]. The mode reduction predicts the functional form of the dynamics governing the resolved, slow variable [60, 61, 62, 63]. The resulting stochastic equation is
$dx=(\tilde{F}+ax+bx^{2}-cx^{3})\,dt+(A-Bx)\,dW_{2}+\sigma\,dW_{3},$ (17)
where the coefficients are expressed in terms of parameters of the triad system in Eq. (6) as: $a=L_{11}+\epsilon(\frac{I^{2}\sigma_{2}^{2}}{2\gamma^{2}_{2}}-\frac{L_{12}^{2}}{\gamma_{2}}-\frac{L_{13}^{2}}{\gamma_{3}})$; $b=-2\epsilon\frac{L_{12}I}{\gamma_{2}}$; $c=\epsilon\frac{I^{2}}{\gamma_{2}}$; $B=-I\epsilon^{1/2}\frac{\sigma_{2}}{\gamma_{2}}$; $L=-B\frac{L_{12}}{I}$; $\tilde{F}=-\frac{AB}{2}$; $\sigma=\epsilon^{1/2}L_{13}\frac{\sigma_{3}}{\gamma_{3}}$. $W_{2}$ and $W_{3}$ are independent Wiener processes. Interestingly, the equation for the coarse-grained variables $x_{1}$ display a structure richer than that of the full triad system in Eq. (6). In addition to the original linear and quadratic interactions, the reduced equation includes new terms: an effective forcing $\tilde{F}$, a cubic nonlinearity, and both additive and multiplicative noise. The additive noise results from the linear interactions across fast modes, while the multiplicative noise reflects the advection of the slow mode by the fast modes [67]. Thus, Eq. (17) demonstrates that multiplicative noise arises from theoretical considerations, rather than being introduced ad hoc, providing a firm basis for its use in reduced-order stochastic climate models. The scalar model above has served as a prototype in several contexts: it has been used to test the Fluctuation–Dissipation Theorem in climate dynamics [81, 114] as well as for inverse modeling of the leading principal component of a quasi-geostrophic model and the observed variability associated with the North Atlantic Oscillation (NAO), demonstrating its relevance for theoretical studies in climate dynamics.
Figure 9: In each panel: $x_{1}$ Triad” represent statistics of $x_{1}$ in the original triad model in Eq. (6); $x_{1}$ Scalar” represent statistics of $x_{1}$ in the scalar reduced model in Eq. (17). Panel (a): stationary probability distribution. Panel (b): response in ensemble mean. Panel (c): response in ensemble variance. Impulse response functions are computed using an ensemble of $N_{e}=10^{7}$ members.
We evaluate the model in Eq. (17) as done in Section III, so to reproduce the stationary statistics of the slow variable $x_{1}$ as well as its response to perturbations as in the original triad system in Eq. (6). The results are reported in Figure 9. The analysis show that, given $\epsilon\ll 1$, it is possible to build a Markovian model of the slow variable $x_{1}$ reproducing both stationary and perturbed statistics as in the original system.
Finally, as shown in this Section, rigorous model reduction strategies such as those developed by Majda and collaborators [61] and by Roberts [134, 135] (among others [2]) offer invaluable theoretical guidelines for multiscale climate modeling. While such strategies where traiditionally developed in the case when the equation of motions are known, they can possibly serve as inspiration and leading to new directions in future work involving data-driven emulators.
Finally, as shown in this Section, rigorous model reduction frameworks, such as those developed by Majda and collaborators [61] and by Roberts [134, 135], among others [2], provide invaluable theoretical guidance for multiscale climate modeling. While these strategies were traditionally developed for systems with known equations of motion, they may also inspire future work aimed at incorporating such principles into data-driven-only emulators.
Appendix C Technical and practical details on the model formulations
C.1 Multivariate data-driven stochastic models